Harvey Habit has a utility function U(c1, c2) = min{c1, c2}, where c1 and c2 are his consumption in periods 1 and 2 respective ly. Harvey earns $189 in period 1 and he will earn $63 in period 2. Harvey can borrow or lend at an interest rate of 10%. There is no inflation.
1
Expert's answer
2016-12-14T11:37:11-0500
1.10*c1 + c2 =((1.10*189)+63) = 270.9 since c1 = c2 at the optimum, we have 1.1*c1+c2 => (1.1*c1+c2)/2.1 =270.9/2.1= → c1 = 129. Then Harvey save 60 because 189-129=60.